OpenStudy (anonymous):

Ellen picks a number. She adds 3 to it and then multiplies the result by 2. She then subtracts 7 from this result. Her final result is 53. If x is the number she originally selected, pick the algebraic equation that represents these operations. 53 = 2(x + 3) - 7 53 = 2x + 3 - 7 7 = 2(x + 3) + 53 53 = 3(x + 2) - 7

OpenStudy (anonymous):

A) \[53 = 2(x + 3) - 7\]

OpenStudy (anonymous):

Translate the following statement into an equation with one variable. The sum of two consecutive even integers is 26. x + y = 26 x + 2x = 26 x + (x + 2) = 26 x + (x + 1) = 26

OpenStudy (anonymous):

Sorry I have to go, if i helped please reward me by giving me a medal when I helped. Thank You!

OpenStudy (anonymous):

it's d: x+(x+1)=26

OpenStudy (anonymous):

\[2(x+3) - 7 =53\]